Given the information in the table, Current dividend $12.00 Growth Rate in Dividends 2% Required Return on Equity Rs 6% According to the Gordon Growth Model, what is the price of this stock in year 2 ?
Answer #1

Po - Divi -9 Po = Price of Stock Divi = Estimated Dividends for Next Period r = Required Rate of Return 9 = Growth Rate

Div3 P21-9

Div3 = Div0 * (1 + g)3 = $12 * (1 + 2%)3 = $12 * 1.0612 = $12.7345

12.7345 = n06 -0.02

P = 318.36